Theres a triangle with angle A=61° angle B=29° and angle C=90°
Side AC=16 and have to find side AB and AC.

Answer by Alan3354(69443)

Theres a triangle with angle A=61° angle B=29° and angle C=90°
Side AC=16 and have to find side AB and AC.
===============
Use AC = b = side opposite angle B
AB = c
BC = a
--------------
sin(B) = b/c = 16/c
c = 16/sin(B)
-------
tan(B) = b/a
a = 16/tan(B)